Consider the function f(x)=(x−2)^3

(a) Estimate the area between the curve y=f(x), the x-axis, and the lines x=2 and x=6, using a Riemann sum with four rectangles (use the right endpoint rule).

(b) Calculate the precise area by taking the limit of a Riemann sum.


1
Expert's answer
2021-11-08T20:33:27-0500

"\newcommand{\disp}{\displaystyle}
\disp
f(x )= (x-2)^3\\
\text{Our change in x($\Delta$) is given by}\\
\Delta x= \frac{6-2}{4}=1\\
\text{The Riemann sum which we denote by $S_n$ is}\\
S_n = \Delta x[f(2+1) +f(2+2) + f(2+3)+f(2+4)]\\
S_n = 1[f(3) +f(4) +f(5) +f(6)]\\
= (3-2)^3+(4-2)^3+(5-2)^3+(6-2)^3=100\\
\text{Suppose [2,6] is divided into n intervals with s the width of each}\\
\text{interval $\frac{6-2}{n} = \frac{4}{n}$}\\
\text{The area of each sum is $\frac{4}{n}(x_i-2)^3$, i = 1, 2, ..., n and}\\
x_i = \frac{4i}{n}\\
\text{Since we are using the right Riemann sum, we have}\\
S = \sum^n_{i=1}\frac{4}{n}(\frac{4i}{n}-2)^3\\
= \frac{1}{n^4}(\sum^n_{i=1} 64i^3-96i^2n+81n^2-8n^3)\\
= \frac{1}{n^4}(64(\frac{n(n+1)}{2})^2-96n(\frac{n(n+1)(2n+1)}{6})+8n^2(\frac{n(n+1)}{2})-8n^4)\\
\text{Taking the limit of the expression above, we have that}\\
\lim_{n \to \infty} \frac{1}{n^4}(64(\frac{n(n+1)}{2})^2-96n(\frac{n(n+1)(2n+1)}{6})+8n^2(\frac{n(n+1)}{2})-8n^4)=12"